OpenStudy (anonymous):

Brad wants to buy flowers for his friend with $15. The daisies are $1 each and the roses are $3 each. He buys 3 more daisies than roses.How much did the roses cost?

OpenStudy (anonymous):

Brad wants to buy flowers for his friend with $15. The daisies are $1 each and the roses are $3 each. He buys 3 more daisies than roses.How much did the roses cost?

OpenStudy (anonymous):

d + 3r = 15 d = r + 3 now sub r + 3 in for din the 1st equation (r + 3) + 3r = 15 4r + 3 = 15 4r = 15 - 3 4r = 12 r = 3 d = r + 3 d = 3 + 3 d = 6 check... d + 3r = 15 6 + 3(3) = 15 6 + 9 = 15 15 = 15 (correct) Brad paid (3 x 3) = $9.00 for the roses (he bought 3 roses) he paid (6 x 1) = $6.00 for the daises (he bought 6 daisies) any questions ?

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

Mitch is buying candy bars for his friend. He wants to give 2 bars to each friend, and he wants to have 10 spare bars. He can afford to buy 28 candy bars. What is the algebraic sentence that will help him figure out how many friends he can treat? F=The number of friends. A: 2f+10=28 B:2f-10=28 C:1/2f-10=28 D:1/2f+10=28